OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. y = 17x2 - 21

OpenStudy (campbell_st):

in this question you have a = 17, b = 0 and c = -21 the formula is \[x=\frac{-b \pm \sqrt{b^4 - 4ac}}{2a}\] substitute the values and then evaluate to find the zeros

OpenStudy (anonymous):

I have tried and always get the wrong answer

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

1.6 Not sure what im doing wrong

OpenStudy (campbell_st):

after substituting you should get \[x =\frac{ 0 \pm \sqrt{0^2 - 4\times17\times-21}}{2\times17}\] this simplifies to \[x =\frac{ \pm \sqrt{1428}}{34}\] which can be simplifed

OpenStudy (anonymous):

btw, that should be a b^2 in that radical in your first post...

OpenStudy (campbell_st):

oops... ty

OpenStudy (anonymous):

didn't matter here anyway... :)

OpenStudy (anonymous):

would it be 1.1?

OpenStudy (campbell_st):

yes.... thats the approximation

OpenStudy (anonymous):

Awesome!